A certain shade of light blue paint,b, is made by mixing 2 1/2 quarts of blue paint with 5 quarts of white paint.

What is the unit rate for 1 cup of blue paint?


1
Expert's answer
2020-12-24T15:05:57-0500

we can write 2 1/2 = 2.5

We can make a proportion:

2.5/5 = 1/x

2.5x = 5

x = 5/2.5

x = 2

2 is the unit rate for 1 cup of blue paint